Freq of autosomal recesive lethal gene is 0.4.find d freq of carrier in a population of 200 individuals.....Ans: 96

Grade:12

1 Answers

Arun
5 years ago
The frequency of the recessive allele is q = 0.4 (given)
 Hence, p = 1-0.4 = 0.6  (as p = 1- q)
According to Hardy- Weinberg principle,  p2 + ?2pq + q2 = 1.
 hnce,  p2 = (0.6)2 => 0.36
                  q2 = (0.4)2 => 0.16
               2pq = 0.48
since, 2pq is the frequency of hetrozygous carrier,
= 0.48 X 200
=> 96 Individuals